OpenStudy (jane11509):

Michael and Susan have a combined age of 17 years. Michael is 5 years older than Susan. How old is Michael? A. 5 B. 6 C. 11 D. 22

OpenStudy (debbieg):

Michael is 5 years older than Susan. If M = Michael's age and S = Susan's age, then how can you express the relationship between M and S algebraically?

OpenStudy (debbieg):

Michael and Susan have a combined age of 17 years. How can you say that algebraically?

OpenStudy (debbieg):

Now use those 2 equations to solve for Michael's age. Solve one of them for S, then substitute that expression into the other equation. That will give you one equation in one unknown (M), solve that, and you have your answer.

OpenStudy (jane11509):

I'm sorry but can you explain that easier please?

OpenStudy (debbieg):

OK. Michael is 5 years older than Susan. Let M be michael's age. Then susan's age = ?? Use the variable M to write an expression for susan's age, given that Michael is 5 years older than Susan.

OpenStudy (jane11509):

Okay

OpenStudy (jane11509):

But how would I write the expression?

OpenStudy (debbieg):

S = ............? Use M in the expression on the right hand side.

OpenStudy (debbieg):

If I said, "Debbie is 20 years older than Jane", then I could write that as: D = J + 20 or alternatively, J = D - 20 See what I mean?

OpenStudy (debbieg):

Michael is 5 years older than Susan. Let M be michael's age. Then S = ......?

OpenStudy (jane11509):

Oh now I see

OpenStudy (jane11509):

thank you ^-^

OpenStudy (jane11509):

So I'd say either Michael was 11 or 22

OpenStudy (jane11509):

Oh so he is 11! xD

OpenStudy (debbieg):

Yup, that's what I get too. Nice job. :)

OpenStudy (jane11509):

Thank you very much :)